OpenStudy (iconforhire56):

A student is trying to solve the system of two equations given below: Equation P: y + z = 6 Equation Q: 3y + 4z = 1 Which of these is a possible step used in eliminating the y-term? (y + z = 6) ⋅ 4 (3y + 4z = 1) ⋅ 4 (y + z = 6) ⋅ −3 (3y + 4z = 1) ⋅ 3

RhondaSommer (rhondasommer):

do you know how to use the elimination theory?

OpenStudy (shadowlegendx):

@IconForHire56 Do you?

OpenStudy (iconforhire56):

not one bit but need a 60 to pass my pre test or i fail the test and won't be able to finish the class

OpenStudy (mathstudent55):

When you use the elimination method, you need to add the two equations to eliminate one variable. For example, if you had the system, of equations: x + y = 8 x - y = 4 you can see that if you add the equations, the y and the -y will add to zero, eliminating the y variable. Then you have a single equation just in x, and you can solve for x.

OpenStudy (mathstudent55):

In your case, if you just add the equations, y + 3y = 4y, and the y is not eliminated. z + 4z = 5z, and the z is not eliminated, so just adding the equations does not work.

OpenStudy (mathstudent55):

In your case, to use the elimination method, you need one extra step. That step is to multiply one of the equations by a number, so that then when you add the equations a variable will be eliminated.

OpenStudy (mathstudent55):

Look at choice A. You multiply the entire first equation by 4. The first equation ends up becoming: 4y + 4z = 24 When you add this equation to the second equation 3y + 4z = 1, 4y + 3y = 7y, and the y is not eliminated, and 4z + 4z = 8z, and the z is not eliminated. That means choice A does not help eliminate a variable, and is not the answer.

OpenStudy (mathstudent55):

Now do the same for the other three choices to see which one will help eliminate a variable.